# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ Abstract
Large language models (LLMs) hold great promise in enhancing psychiatric practice, from improving diagnostic accuracy to streamlining clinical documentation and therapeutic support. However, existing evaluation resources heavily rely on small clinical interview corpora, social media posts, or synthetic dialogues, which limits their clinical validity and fails to capture the full complexity of psychiatric reasoning. In this work, we introduce PsychiatryBench, a rigorously curated benchmark grounded exclusively in authoritative, expert-validated psychiatric textbooks and casebooks (e.g., DSM-5-TR Clinical Cases, Stahl’s Essential Psychopharmacology). PsychiatryBench comprises eleven distinct question-answering tasks ranging from diagnostic reasoning and treatment planning to longitudinal follow-up, management planning, clinical approach, sequential case analysis, and multiple-choice/extended matching formats totaling over 5,300 expert-annotated items. We evaluate a diverse set of frontier LLMs (including Google Gemini, DeepSeek, LLaMA 3, and QWQ-32) alongside leading open-source medical models (e.g., OpenBiloLLM, MedGemma) using both conventional metrics and an "LLM-as-judge" similarity scoring framework. Our results reveal substantial gaps in clinical consistency and safety, particularly in multi-turn follow-up and management tasks, underscoring the need for specialized model tuning and more robust evaluation paradigms. PsychiatryBench offers a modular, extensible platform for benchmarking and improving LLM performance in high-stakes mental health applications.
@@ 1 Introduction
Large language models (LLMs), a subset of artificial intelligence (AI), have the potential to completely transform the way mental illness is identified, diagnosed, and treated. In 2019, mental disorders accounted for 5.1% of the world’s disease burden, with over 280 million people suffering from depressive disorders and around 703,000 people dying from suicide [1]. Between 2024 and 2040, the total economic burden of mental health problems in the United States is expected to reach around \$14 trillion, including direct medical expenses, emergency room visits, lost productivity, and premature death [2].
New avenues for mental health research were made possible with the introduction of Natural Language Processing (NLP). Early applications focused on analyzing text from self-report questionnaires and clinical interviews using techniques like the Linguistic Inquiry and Word Count (LIWC) [3]. A significant advancement came with transformer-based models such as BERT [4], which learn complex linguistic representations from large text corpora without manual feature engineering. These models have shown strong performance in tasks like sentiment analysis, suicide risk assessment, and detecting mental health disorders from social media content [5].
However, despite these advances, there remains a critical gap: existing evaluation frameworks rarely test models on tasks that reflect the depth and complexity of real-world psychiatric decision-making. Most benchmarks rely on simplified labels or non-clinical data, limiting their relevance to actual clinical practice. To address this, our study introduces a new benchmark grounded in expert-curated psychiatric sources and designed to assess nuanced reasoning tasks such as differential diagnosis and treatment planning. This framework not only raises the bar for clinical relevance in LLM evaluation but also enables robust comparison across state-of-the-art models.
The integration of LLMs into mental healthcare represents a significant technological shift, with the potential to transform clinical practice across the entire psychiatric workflow[6]. In diagnostics, LLMs demonstrate effectiveness in detecting disorders like depression from textual data; in therapeutic interventions, through conversational agents; and in clinical documentation, their automated notes have been preferred by therapists in blind tests over human-written ones[7]. Intriguingly, some studies suggest that LLMs can even surpass human physicians in expressing empathy in written communication[8]. This transformative potential, however, is counterbalanced by significant risks that necessitate an emerging field that can be termed ’Evaluation Psychiatry.’ A primary concern is the propensity of LLMs to generate inaccurate or nonsensical responses, known as "hallucinations," which raises substantial safety concerns in a clinical domain where errors can have profound consequences pandey2024harnessinglargelanguagemodels. This tension between the promise of democratizing mental health access and the ethical imperative of patient safety underscores the critical need for rigorous, clinically-grounded evaluation frameworks.
The root of this tension can be traced to the evolution of the datasets used to evaluate these models. Early, influential resources like the Distress Analysis Interview Corpus (DAIC-WOZ) provided clinically relevant transcripts that enabled foundational methods, but these datasets were small and faced challenges in generalizability [9]. To achieve greater scale, a second wave of datasets leveraged social media and public surveys, such as the subreddit Mental Health Posts dataset[10],[11]. Models like MentalBERT, pretrained on millions of sentences from mental health subreddits, achieved high accuracy (over 94%) on simplified tasks like binary depression classification[12]. While valuable for demonstrating potential, this approach has a critical flaw: its reliance on unverified, non-clinical data oversimplifies complex psychiatric disorders and risks reflecting trends in self-diagnosis rather than formal clinical presentations.
More recently, the field has begun to address this by developing more complex, conversational benchmarks. MentalChat16K, for example, combines synthetic data with anonymized clinical transcripts to evaluate an LLM’s ability to engage in nuanced dialogue[13]. Reflects a maturation of the field, moving beyond simple classification. Nonetheless, this third wave introduces new methodological challenges. The reliance on synthetic data risks creating benchmarks that do not capture the full complexity of real human interaction. At the same time, the use of LLM-as-a-judge evaluation methods introduces a circular dependency, where the technology being assessed is also the tool for its assessment, potentially masking its own biases and unreliability.
@@
This analysis of the current landscape reveals a critical and persistent gap: the lack of a comprehensive benchmark that is simultaneously grounded in authoritative, expert-validated knowledge and designed to evaluate the full spectrum of complex clinical reasoning. There is a stark contrast between the datasets currently in use and the foundational knowledge sources that form the bedrock of clinical psychiatric practice. Authoritative textbooks and casebooks, such as the DSM-5-TR Clinical Cases[14] and Stahl’s Essential Psychopharmacology[15], represent knowledge curated and validated by experts over decades. These resources, which illustrate the complexities of differential diagnosis and treatment planning, stand as the gold standard against which an LLM’s clinical acumen and safety must ultimately be measured.
The present study is positioned to address these limitations directly. We introduce a novel evaluation framework meticulously constructed from the foundational psychiatric textbooks described above, deliberately prioritizing clinical accuracy, safety, and alignment with established medical practice. Our contribution is threefold:
@@
  • We construct a novel, diverse dataset curated from authoritative clinical casebooks, diagnostic manuals, and specialized guides, ensuring that our evaluation is grounded in validated, expert knowledge rather than unverified or synthetic data.
  • We design eleven distinct and complex QA task types, ranging from diagnostic reasoning and treatment planning to sequential, case-based analysis, to create a framework that evaluates the sophisticated cognitive functions required in real-world psychiatry, moving far beyond simple classification.
  • We deploy this framework to conduct a rigorous evaluation of leading frontier models, establishing a crucial benchmark that provides novel insights into the current capabilities and inherent risks of LLMs in a clinical context.
  • A comparative evaluation of several state-of-the-art LLMs, including Gemini 2.5 Pro, Gemini 2.5 flash, Claude Opus, and Qwen/QWQ-32, selected for their strong performance across reasoning and safety benchmarks, and for their relevance in current clinical LLM discourse.
@@
This work, therefore, moves beyond testing what LLMs can do on simplified tasks and begins to assess their readiness for the complex, rigorous, high-stakes reality of clinical mental healthcare. The remainder of the paper is structured as follows: Section 2 provides a detailed review of LLM applications in mental health and the landscape of psychiatric datasets. Section 3 describes the design, sourcing, and annotation of our novel psychiatric QA dataset. Section 4 outlines the experimental framework, and Section 5 details the implementation. Section 6 presents and interprets our results. Section 7 discusses study limitations and future work, and Section 8 concludes the paper by summarizing our key contributions. Finally, Ethical, Copyright, and Fair Use Statements are presented.
@@ 2.1 The Emergence of LLMs in Mental Healthcare
The integration of LLMs into healthcare represents a significant technological shift, with the potential to transform clinical practice, particularly in mental health [16]. The field is witnessing a rapid expansion of LLM applications across the entire psychiatric workflow. In diagnostics, LLMs demonstrate effectiveness in detecting mental health disorders and suicidal ideation from textual data, often sourced from social media, and they can translate complex diagnostic manuals like the ICD-11 into inspectable logic programs for expert review [17]. For therapeutic interventions, conversational agents such as Woebot and Wysa, built on principles of cognitive behavioral therapy, are providing accessible and de-stigmatized eHealth services. Intriguingly, some studies suggest that LLMs can even surpass human physicians in expressing empathy in written communication [18]. Furthermore, in clinical documentation, LLMs offer dramatic efficiency gains by automating the generation of medical notes. In this task, therapists have preferred their output in blind tests over human-written notes [19].
This transformative potential, however, is counterbalanced by significant risks that necessitate an emerging field that can be termed ’Evaluation Psychiatry.’ A primary concern is the propensity of LLMs to generate inaccurate or nonsensical responses, known as "hallucinations," which raises substantial safety concerns in a clinical domain where errors can have profound consequences [20]. This risk mandates stringent ethical guidelines and a "human-in-the-loop" paradigm where AI outputs are subject to expert validation. Clinicians also express ambivalence, worrying about over-reliance on "black box" algorithms, the erosion of clinical autonomy, and the potential for training data to embed biases related to socioeconomic status or geography, which could exacerbate healthcare disparities [21]. This tension between the promise of democratizing mental health access and the ethical imperative of patient safety underscores the critical need for rigorous, clinically-grounded evaluation frameworks.
@@ 2.2 The Landscape of Datasets for Psychiatric Evaluation
The evolution of LLM evaluation in mental health can be traced through the datasets used to train and test them. Early and influential resources include clinical interview corpora like the Distress Analysis Interview Corpus (DAIC-WOZ) and its extension, E-DAIC.[22] These datasets, containing transcripts of semi-structured interviews and corresponding psychiatric assessment scores (e.g., PHQ-8 for depression), have been instrumental. They enabled innovative methods such as the Language Model for Impersonation-based Questionnaire Completion (LMIQ) framework, which attempts to bridge unstructured patient narratives with quantitative metrics [18]. While a significant step, these methods face challenges in generalization beyond their specific benchmarks.
A second wave of datasets has leveraged the scale of social media and public surveys. Resources like the Reddit Mental Health Posts dataset, the Depression Detection Using Text corpus, and various mental health surveys have provided vast amounts of natural language data for tasks like sentiment analysis and binary classification [23]. Models like Mental BERT and Mental RoBERTa were pretrained on millions of sentences from mental health-related subreddits to improve performance in these tasks [17]. However, the clinical validity of social media data is a significant concern due to its unverified nature, oversimplification of symptoms, and the risk of reflecting trends in self-diagnosis rather than formal clinical presentations.
More recently, the field has begun to address the need for more complex, conversational benchmarks. MentalChat16K, for instance, is a benchmark for conversational mental health assistance that combines synthetic data with anonymized clinical transcripts. [13]. Simultaneously, researchers have also investigated automated approaches for enhancing dataset richness and annotation quality. For example, Hassan et al.(2024) introduced a synthetic labeling framework that transforms traditional single-disorder datasets into multi-label corpora by leveraging zero-shot and prompt-driven LLM annotation strategies. Their method, exemplified by the SPAADE-DR dataset, better captures co-occurring disorders such as depression, anxiety, PTSD, ADHD, and eating disorders, thus enabling more nuanced diagnostic modeling[24]. This reflects a maturation of the field, moving beyond simple classification towards evaluating LLMs on their ability to engage in nuanced, therapeutic dialogue. Nonetheless, the reliance on synthetic data and LLM-as-a-judge evaluation methods introduces new layers of potential bias and unreliability.
@@ 2.3 Foundational Knowledge Sources in Psychiatric Practice
  • Diagnostic Manuals and Casebooks: Books like DSM-5 Clinical Cases and its successor, DSM-5-TR Clinical Cases, are essential companions to the official diagnostic manual. They bring diagnostic criteria to life through realistic narratives and detailed clinical discussions, illustrating the complexities of comorbidity and differential diagnosis. Similarly, texts such as 100 Cases in Psychiatry and Case Files Psychiatry provide high-yield clinical scenarios designed to sharpen practical decision-making skills, making them ideal for evaluating an LLM’s clinical reasoning.
  • Specialized Clinical Guides: Deep domain expertise is captured in specialized resources. Stahl’s Essential Psychopharmacology and its case-based companions are seminal texts on the neurobiological basis and practical application of psychiatric medications, crucial for evaluating treatment-related tasks. Likewise, Geriatric Psychiatry provides comprehensive knowledge on the unique diagnostic and treatment considerations for elderly populations.
  • Structured Knowledge and Assessment Tools: To test factual recall and application of diagnostic criteria in a standardized format, resources like DSM-5-TR Self-Exam Questions and MCQs in Psychology are invaluable. They provide thousands of multiple-choice questions that mirror the format of professional board examinations, offering a structured way to measure an LLM’s knowledge base.

### v2

# PsychiatryBench: A Multi-Task Benchmark for LLMs in Psychiatry

@@ Abstract
Large language models (LLMs) offer significant potential in enhancing psychiatric practice, from improving diagnostic accuracy to streamlining clinical documentation and therapeutic support. However, existing evaluation resources heavily rely on small clinical interview corpora, social media posts, or synthetic dialogues, which limits their clinical validity and fails to capture the full complexity of diagnostic reasoning. In this work, we introduce PsychiatryBench, a rigorously curated benchmark grounded exclusively in authoritative, expert-validated psychiatric textbooks and casebooks. PsychiatryBench comprises eleven distinct question-answering tasks ranging from diagnostic reasoning and treatment planning to longitudinal follow-up, management planning, clinical approach, sequential case analysis, and multiple-choice/extended matching formats totaling 5,188 expert-annotated items. We evaluate a diverse set of frontier LLMs (including Google Gemini, DeepSeek, Sonnet 4.5, and GPT 5) alongside leading open-source medical models such as MedGemma using both conventional metrics and an "LLM-as-judge" similarity scoring framework. Our results reveal substantial gaps in clinical consistency and safety, particularly in multi-turn follow-up and management tasks, underscoring the need for specialized model tuning and more robust evaluation paradigms. PsychiatryBench offers a modular, extensible platform for benchmarking and improving LLM performance in mental health applications.
@@ 1 Introduction
Mental disorders impose a staggering global burden, accounting for 5.1% of worldwide disease burden in 2019, affecting over 280 million people with depression, and claiming 703,000 lives to suicide [WHO2019]. The economic toll is projected to reach \$14 trillion in the United States alone between 2024 and 2040, encompassing direct medical costs, emergency interventions, productivity losses, and premature mortality [Smith2023]. These figures demand urgent action: scalable technologies that can improve diagnostic precision, expand care access, and accelerate treatment delivery are not merely aspirational; they are critical infrastructure gaps in modern healthcare systems.
Large Language Models (LLMs), a subset of artificial intelligence (AI), have emerged as a potentially transformative technology in healthcare. The integration of LLMs into psychiatric practice represents a significant technological development with substantive implications for clinical practice. Recent studies have demonstrated that models can perform meaningful tasks in mental health contexts, including detecting depression and suicide risk from text, providing supportive information through conversational agents, and generating clinical documentation that psychiatrists prefer in blind comparisons [shah2025tnevalrubricevaluationprotocols, ji2021mentalbert]. Some investigations have even suggested that LLMs can surpass human physicians in expressing empathy in written clinical communication [Hua2025LLMsMentalHealth]. These capabilities suggest that LLMs may offer pathways to democratize mental health assessment and support, potentially reaching populations with limited access to specialized psychiatric care.
However, the transformative potential of LLMs in psychiatry is fundamentally constrained by substantial safety and reliability concerns. A primary concern is the well-documented propensity of LLMs to generate inaccurate or nonsensical responses, a phenomenon known as “hallucinations,” which raises profound safety issues in psychiatry where diagnostic errors can result in serious patient harm [pandey2024harnessinglargelanguagemodels]. The landscape of existing evaluation frameworks reveals critical limitations. Early, influential resources provided clinically relevant data but were limited in scale and faced significant challenges in generalizability [Guo_2024]. To achieve broader evaluation scope, subsequent benchmarks increasingly relied on unverified social media data, which tends to oversimplify complex psychiatric presentations and risks conflating self-identified symptoms with formal clinical diagnoses rather than representing the diagnostic reasoning required in professional psychiatric practice [low2020natural, hanafi2024comprehensiveevaluationlargelanguage]. More recent efforts have developed conversational benchmarks addressing clinical dialogue, but these frequently depend on synthetic data or employ other LLMs as evaluators, introducing methodological concerns and potential biases that compromise validity [xu2025mentalchat16kbenchmarkdatasetconversational, li2024llmsasjudgescomprehensivesurveyllmbased].
Despite these advances, a critical evaluation gap persists: existing frameworks rarely assess models on tasks that genuinely reflect the depth, complexity, and rigor of real-world psychiatric clinical reasoning and decision-making. This gap is particularly consequential because psychiatry, as a medical specialty, demands the integration of diagnostic criteria, systematic differential diagnosis, treatment algorithm selection, and dynamic case management competencies that go far beyond pattern recognition or conversational fluency.
@@
The authoritative knowledge foundations of psychiatric practice represented in curated clinical casebooks and diagnostic resources such as the DSM-5-TR Clinical Cases [APA2013] and Stahl’s Essential Psychopharmacology [stahl2021essential] represent decades of expert knowledge curation and validation. These resources document the complexities of differential diagnosis, the nuances of treatment planning, and the sophisticated clinical judgment required in real psychiatric practice. These foundational sources establish the gold standard against which an LLM’s clinical reasoning, diagnostic acumen, safety awareness, and appropriateness for psychiatric applications must ultimately be measured[aljohani2025comprehensivesurveytrustworthinesslarge].
To address this critical evaluation gap, we introduce PsychiatryBench, a novel evaluation framework meticulously constructed from foundational psychiatric textbooks and expert-validated clinical cases, deliberately prioritizing clinical accuracy, safety, and alignment with established medical practice. PsychiatryBench is designed with a single, exclusive purpose: to serve as a rigorous, standardized tool for comparative evaluation of LLM’s capabilities in psychiatry. Its core aim is to provide a benchmark that assesses whether LLMs can approximate expert-level psychiatric reasoning and clinical decision-making as demonstrated by experienced psychiatrists.
We define high-stakes psychiatric applications as clinical contexts where AI errors could result in serious patient harm, including: misdiagnosis leading to inappropriate or delayed treatment, failure to recognize suicide risk, incorrect medication recommendations causing adverse effects, or inappropriate crisis management. Such applications demand rigorous evaluation standards that assess not only general language understanding but also clinical accuracy, and appropriate acknowledgment of model uncertainty and limitations.
Scope: PsychiatryBench is intentionally focused on adult psychiatry as a medical specialty within the broader mental health field. The benchmark targets psychiatry-specific clinical reasoning and decision-making, excluding general mental health interventions, psychotherapy, counseling techniques, and broader behavioral health domains. Child and adolescent psychiatry was deliberately excluded, as it represents a distinct subspecialty with unique developmental, diagnostic, and therapeutic considerations that would necessitate a dedicated evaluation framework. The inclusion of geriatric psychiatry reflects the clinical reality that geriatric cases are commonly encountered by general adult psychiatrists. While source cases include references to medical mimics of psychiatric conditions (e.g., thyroid disease, CNS pathology, delirium), these are not systematically or comprehensively represented across all tasks. The benchmark does not systematically assess conversational coherence (turn-by-turn dialogue management) or narrative biopsychosocial formulation (holistic case conceptualization integrating biological, psychological, and social factors), as our focus is on knowledge application and clinical decision-making rather than interactive communication or narrative synthesis.
Throughout this work, the term psychiatric reasoning is used as an umbrella term encompassing six specific, complementary dimensions of clinical competence, each targeted through dedicated task types:
Diagnostic reasoning: The cognitive process of synthesizing patient presentation, history, and mental status examination findings to generate and systematically differentiate among possible psychiatric disorders (assessed through Diagnosis and Classification tasks)
Treatment decision-making: Clinical judgment in selecting appropriate therapeutic interventions based on diagnosis, patient characteristics, individual vulnerabilities, and evidence-based guidelines (assessed through Treatment and Treatment Follow-Up tasks)
Management planning: Comprehensive care coordination integrating diagnostic formulation, risk assessment, therapeutic interventions, psychoeducation, and coordinated follow-up strategies (assessed through Management Plan tasks)
Clinical approach: The structured reasoning process for systematically evaluating clinical presentations, generating diagnostic hypotheses, and prioritizing clinical investigations (assessed through Clinical Approach tasks)
Foundational knowledge: Factual recall and application of psychiatric concepts, diagnostic criteria, treatment mechanisms, and clinical principles (assessed through Mental QA, MCQ, and EMI tasks)
Sequential reasoning: The ability to track evolving clinical information across multiple time points and adapt diagnostic and treatment recommendations accordingly (assessed through Sequential QA tasks).
Our contribution comprises three interconnected elements:
@@
  • We construct a novel, clinically rigorous dataset curated from authoritative psychiatric casebooks, diagnostic manuals, and specialized clinical guides, ensuring that evaluation is grounded in validated, expert knowledge rather than unverified or synthetically generated data.
  • We design eleven distinct and complex task types that move beyond simplified assessment to rigorously evaluate LLMs’ readiness for the complex, intellectually demanding, and ethically high-stakes reality of contemporary psychiatric clinical practice.
  • We provide a standardized evaluation resource specifically designed for comparative LLM assessment in psychiatry. While insights from this benchmark may secondarily inform discussions about clinical decision support systems or educational applications, PsychiatryBench’s primary and exclusive purpose is rigorous model evaluation and comparison. We acknowledge that application to regulatory or policy decisions would necessitate substantial expansion to address cultural, developmental, legal, and systems-level domains, which are explicitly outside the scope of this work.
@@
The remainder of this paper is structured as follows: Section 2 provides a detailed review of LLM applications in psychiatry and the landscape of existing psychiatric evaluation datasets. Section 3 describes the design, sourcing, and annotation methodology for our novel psychiatric evaluation dataset. Section 4 outlines the experimental design and evaluation methodology, including model implementation details and evaluation protocols. Section 5 presents and interprets our empirical results. Section 6 discusses study limitations and directions for future work. Section 7 concludes by summarizing our key contributions and implications. Ethical, Copyright, and Fair Use Statements are presented in section 8.
@@ 2.1 The Emergence of LLMs in Mental Healthcare
The integration of LLMs into healthcare represents a significant technological shift, with the potential to transform clinical practice particularly within mental health care by leveraging their remarkable capabilities in language understanding and generation [Sahu2025, LIN2025100868]. The field is witnessing a rapid expansion of LLM applications across the entire psychiatric workflow. In diagnostics, LLMs demonstrate effectiveness in detecting mental health disorders and suicidal ideation from textual data, often sourced from social media, and they can translate complex diagnostic manuals such as the ICD-11 into inspectable logic programs suitable for expert review [na2025surveylargelanguagemodels].
For therapeutic interventions, conversational agents such as Woebot and Wysa, grounded in principles of cognitive behavioral therapy, offer accessible and de-stigmatized eHealth services [Fitzpatrick2017, Beatty2022]. Emerging research even from non-psychiatric contexts like primary care suggests that LLMs may in some cases be perceived as more empathetic or higher-quality communicators than human clinicians in written exchanges [rousmaniere2025large, Hatch2025, Kuhail2024]. In clinical documentation, LLMs have demonstrated substantial potential to improve efficiency by automating the generation of therapy and medical notes. Notably, therapists have preferred LLM-generated behavioral health notes over human-written ones in blind assessments [StanfordHAI2025, shah2025tnevalrubricevaluationprotocols].
A primary concern is the propensity of LLMs to generate inaccurate or nonsensical outputs commonly referred to as “hallucinations” which poses substantial safety concerns in clinical settings where errors carry serious consequences [aljohani2025comprehensivesurveytrustworthinesslarge]. These issues necessitate stringent ethical guidelines and reinforce the importance of maintaining a “human-in-the-loop” framework in which clinician oversight remains essential.
Clinicians also express ambivalence about widespread LLM adoption, citing worries about over-reliance on opaque “black-box” algorithms, potential diminution of clinical autonomy, and biases embedded in training data related to socioeconomic status or geography. Such biases risk exacerbating existing healthcare disparities [Mancini2025]. This tension between the promise of democratizing access to mental health support and the ethical imperative of ensuring patient safety highlights the urgent need for rigorous, clinically grounded evaluation frameworks that can guide the responsible integration of LLMs into psychiatric practice.
@@ 2.2 The Landscape of Datasets for Psychiatric Evaluation
The evolution of LLM evaluation in mental health can be traced through the datasets used to train and test them. Early and influential resources include clinical interview corpora like the Distress Analysis Interview Corpus (DAIC-WOZ) and its extension, E-DAIC.[ali2024leveragingaudiotextmodalities] These datasets, containing transcripts of semi-structured interviews and corresponding psychiatric assessment scores (e.g., PHQ-8 for depression), have been instrumental. They enabled innovative methods such as the Language Model for Impersonation-based Questionnaire Completion (LMIQ) framework, which attempts to bridge unstructured patient narratives with quantitative metrics [rousmaniere2025large]. While a significant step, these methods face challenges in generalization beyond their specific benchmarks.
A second wave of datasets has leveraged the scale of social media and public surveys. Resources like the Reddit Mental Health Posts dataset, the Depression Detection Using Text corpus, and various mental health surveys have provided vast amounts of natural language data for tasks like sentiment analysis and binary classification [Jin2025]. Models like Mental BERT and Mental RoBERTa were pretrained on millions of sentences from mental health-related subreddits to improve performance in these tasks [na2025surveylargelanguagemodels]. However, the clinical validity of social media data is a significant concern due to its unverified nature, oversimplification of symptoms, and the risk of reflecting trends in self-diagnosis rather than formal clinical presentations.
More recently, the field has begun to address the need for more complex, conversational benchmarks. MentalChat16K, for instance, is a benchmark for conversational mental health assistance that combines synthetic data with anonymized clinical transcripts. [xu2025mentalchat16kbenchmarkdatasetconversational]. Simultaneously, researchers have also investigated automated approaches for enhancing dataset richness and annotation quality. For example, [hassan2024automatedmultilabelannotationmental] introduced a synthetic labeling framework that transforms traditional single-disorder datasets into multi-label corpora by leveraging zero-shot and prompt-driven LLM annotation strategies. Their method, exemplified by the SPAADE-DR dataset, better captures co-occurring disorders such as depression, anxiety, PTSD, ADHD, and eating disorders, thus enabling more nuanced diagnostic modeling. This development reflects the maturation of the field, as research is moving beyond simple classification toward evaluating LLMs on their ability to engage in nuanced, therapeutic dialogue. Nonetheless, the reliance on synthetic data and LLM-as-a-judge evaluation methods introduces new layers of potential bias and unreliability.
@@ 2.3 Foundational Knowledge Sources in Psychiatric Practice
  • Diagnostic Manuals and Casebooks: Books like DSM-5 Clinical Cases[barnhill2013dsm5-clinical-cases] and its successor, DSM-5-TR Clinical Cases[Barnhill2022_DSM5TRClinicalCases], are essential companions to the official diagnostic manual. They bring diagnostic criteria to life through realistic narratives and detailed clinical discussions, illustrating the complexities of comorbidity and differential diagnosis. Similarly, texts such as 100 Cases in Psychiatry[wright2017_100casespsychiatry] and Case Files Psychiatry[toy2009casefiles-psychiatry] provide high-yield clinical scenarios designed to sharpen practical decision-making skills, making them ideal for evaluating an LLM’s clinical reasoning.
  • Specialized Clinical Guides: Deep domain expertise is captured in specialized resources. Stahl’s Essential Psychopharmacology[stahl2021essential] and its case-based companions are seminal texts on the neurobiological basis and practical application of psychiatric medications, crucial for evaluating treatment-related tasks. Likewise, Geriatric Psychiatry[hategan2024geriatric] provides comprehensive knowledge on the unique diagnostic and treatment considerations for elderly populations.
  • Structured Knowledge and Assessment Tools: To test factual recall and application of diagnostic criteria in a standardized format, resources like DSM-5-TR Self-Exam Questions are invaluable. They provide thousands of multiple-choice questions that mirror the format of professional board examinations, offering a structured way to measure an LLM’s knowledge base.
@@
The DSM-5-TR (Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition, Text Revision; American Psychiatric Association, 2022) is the current edition. Our source materials include textbooks and casebooks aligned with both the original DSM-5 (2013) and DSM-5-TR. We refer to "DSM-5-TR" when discussing current diagnostic criteria and standards, and specify "DSM-5" only when citing pre-2022 source materials by their original titles.